# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v2

@@
Abstract
While the progression of Large Language Models (LLMs) has notably propelled financial analysis, their application has largely been confined to singular language realms, leaving untapped the potential of bilingual Chinese-English capacity. To bridge this chasm, we introduce ICE-PIXIU111https://github.com/YY0649/ICE-PIXIU, seamlessly amalgamating the ICE-INTENT model and ICE-FLARE benchmark for bilingual financial analysis. ICE-PIXIU uniquely integrates a spectrum of Chinese tasks, alongside translated and original English datasets, enriching the breadth and depth of bilingual financial modeling. It provides unrestricted access to diverse model variants, a substantial compilation of diverse cross-lingual and multi-modal instruction data, and an evaluation benchmark with expert annotations, comprising 10 NLP tasks, 20 bilingual specific tasks, totaling 95k datasets. Our thorough evaluation emphasizes the advantages of incorporating these bilingual datasets, especially in translation tasks and utilizing original English data, enhancing both linguistic flexibility and analytical acuity in financial contexts. Notably, ICE-INTENT distinguishes itself by showcasing significant enhancements over conventional LLMs and existing financial LLMs in bilingual milieus, underscoring the profound impact of robust bilingual data on the accuracy and efficacy of financial NLP.
@@ Abstract
1 Introduction
@@
Inspired by the success of general-domain large language models (LLMs), the exploration of LLMs in the financial sector is gaining momentum Shah et al. (2022); Lu et al. (2023); Xie et al. (2023b, 2024, a). Starting with the broad capabilities of models like GPT-3.5 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), the focus has shifted towards more specialized models such as BloombergGPT (Wu et al., 2023), which tackles the complex nuances of financial language and concepts. Subsequently, models like FinGPT (Yang et al., 2023b), InvestLM (Yang et al., 2023c) and Typhoon (Pipatanakul et al., 2023) emerged, concentrating on English and Thai financial tasks. Notably, FinMA (Xie et al., 2023b) addresses the challenges of instruction data modality within this domain, marking a significant advancement.
@@
Beyond the efforts in English-centric models, Chinese financial LLMs (FinLLMs) have also made notable advances. XuanYuan2.0 (Zhang et al., 2023d) exemplifies this trend, focusing on the intricacies of Chinese financial language. This has been followed by models such as DISC-FinLLM (Chen et al., 2023) and CFGPT (Li et al., 2023a), which further the capabilities in financial analysis and reasoning. Recently, PanGu-π (Wang et al., 2023) enhances a pretrained model for better performance in financial examinations. Moreover, certain business-oriented LLMs (Wang, 2023) like TigerBot and LightGPT remain closed-source.
@@
| Model | Size | Language | Instruction | Evaluation | Open Source | Modality | Release Date | ||||
| zh | en | zh | en | dof | Model | Data | Only Text | ||||
| BloombegGPT (Wu et al., 2023) | 50B | en | 0 | 5 | 0 | 5 | 0 | 03/30/23 | |||
| InvestLM (Yang et al., 2023c) | 65B | en | 0 | 9 | 0 | 9 | 0 | 09/14/23 | |||
| FinGPT (Wu et al., 2023) | 7/13B | en | 0 | 6 | 0 | 6 | 0 | 11/10/23 | |||
| XuanYuan2.0 (Zhang et al., 2023d) | 176B | zh | - | 0 | 0 | 1 | 0 | 05/19/23 | |||
| FinMA (Xie et al., 2023b) | 7/13B | en | 0 | 8 | 0 | 8 | 0 | 06/01/23 | |||
| CFGPT (Li et al., 2023a) | 7B | zh | 6 | 0 | 6 | 0 | 0 | 09/01/23 | |||
| DISC-FinLLM (Chen et al., 2023) | 13B | zh | 8 | 0 | 8 | 0 | 0 | 10/24/23 | |||
| PanGu-π (Wang et al., 2023) | 1/7B | zh | - | 0 | 4 | 0 | 0 | 12/27/23 | |||
| ICE-INTENT | 7B | zh, en | 12 | 6 | 12 | 6 | 3 | 12/10/23 | |||
However, there is a significant challenge in their application and assessment in bilingual contexts, especially regarding model development, instruction dataset diversity, and evaluation methodologies in English and Chinese. As Table 1 illustrates, most FinLLMs are tailored exclusively for either English or Chinese, which underscores a significant shortfall in models proficient in both languages.
This issue is further compounded in the domain of instruction datasets. While there are tasks encompassing classification, extraction, prediction and reasoning, these datasets, particularly for Chinese, show a notable lack of diversity in each task category (Chen et al., 2023; Li et al., 2023a). Furthermore, the original English datasets and their subsequent translations into Chinese, are often overlooked in the development of comprehensive bilingual instruction datasets (Xie et al., 2023b).
Correspondingly, the evaluation methodologies for these models tend to mirror these linguistic limitations (Zhang et al., 2023b; Lei et al., 2023; Zhang et al., 2023c), focusing mainly on monolingual assessments and failing to comprehensively evaluate bilingual proficiency. There is a critical need for more diverse Chinese instruction datasets in both classification, extraction, reasoning and prediction, alongside a greater focus on the translation and utilization of original English datasets into Chinese.
To address this issue, we introduce ICE-PIXIU, a comprehensive framework that features ICE-INTENT and ICE-FLARE, the first cross-lingual bilingual financial LLM and evaluation benchmark, respectively. This framework is characterized by several key attributes:
Bilingual Capability Proficiency: ICE-INTENT, a component of ICE-PIXIU, demonstrates outstanding bilingual proficiency in English and Chinese, essential for global financial data processing and analysis.
Diverse Chinese Datasets: ICE-PIXIU addresses gaps in Chinese financial NLP by incorporating a variety of Chinese classification, extraction, reasoning and prediction NLP tasks, enhancing training and performance.
Professionally Annotated Prompts: ICE-PIXIU offers a a diverse, high-quality, expert-annotated set of prompts applied with similar fine-tuning instructions to enhance understanding of financial tasks.
Incorporation of Translation and English Data: The framework expands its capabilities by including translation tasks and English datasets, bolstering its bilingual training and application.
Cross-Lingual Evaluation with ICE-FLARE: ICE-PIXIU introduces ICE-FLARE, a rigorous cross-lingual evaluation benchmark, ensuring consistent model performance in diverse linguistic contexts.
Open-Source Contribution: With its open-access approach, ICE-PIXIU offers its resources to the research community, fostering collaborative advancement in financial NLP.
@@
In the creation of ICE-INTENT, we meticulously gathered 40 datasets, which consisted of 1,185,076 raw data, 603,940 instruction fine-tuning data, and 95,091 evaluation data, strategically covering a wide range of financial tasks. This ensemble, as detailed in Table 1, comprises 18 specific task on Chinese-English bilingual instruction datasets. This diverse array not only underscores our model’s commitment to linguistic breadth but also to task-specific depth, ensuring ICE-INTENT’s proficiency across various financial scenarios 222Detailed examples are shown in Appendix Figure 8-9. Complementing this, ICE-FLARE, our cross-lingual evaluation benchmark, incorporates 21 evaluation specific tasks—12 in Chinese and 9 in English. Particularly, we innovatively introduced data-out-of-field specific task into ICE-FLARE. This robust framework highlights the unique strength of our approach in providing comprehensive cross-lingual consistency and generalization assessments, crucial for bilingual FinLLMs.
In our evaluation conducted using ICE-FLARE, we observed distinct performance patterns of our financial LLM ICE-INTENT in comparison with advanced general-domain LLMs and existing FinLLMs. The results firstly emphasized the model’s enhanced capabilities in cross-lingual tasks, particularly in handling complex and lengthy text, underscoring the critical role of bilingual data integration in financial language modeling. Subsequently, the model exhibited proficiency in translation tasks, highlighting its adeptness at managing linguistic variations crucial for global financial contexts. Finally, our model outperformed general LLMs in tasks specifically fine-tuned for the financial domain but faced challenges in less familiar tasks within ICE-FLARE. This delineates the specialized nature and unique challenges of the financial sector, reinforcing the need for tailored FinLLMs.
2 Related Work
@@
Financial Dataset Resource: Financial pre-trained language models (FinPLMs) previously focused on single-task data but now most FinLLMs prioritize specific-task data collection. For instance, BloombergGPT (Wu et al., 2023) integrates general and financial text for training large 500B LLMs. FinLLMs like FinGPT (Yang et al., 2023b; Zhang et al., 2023a) curate English instruction datasets for classification, extraction, and generation NLP tasks. PIXIU (Xie et al., 2023b) provide open-source, multi-task and multi-modal data covering diverse financial scenarios. BBT-Fin (Lu et al., 2023) Corpus aggregates Chinese financial data including corporate announcements, research reports, social media, and more, while XuanYuan (轩辕) 2.0 (Zhang et al., 2023d) offers a hybrid-tuning dataset. Cornucopia’s (Yu, 2023) 26M instruction dataset covers application areas like insurance and stocks. Most fine-tuning datasets are tailored for monolingual tasks, limiting coverage and open sourcing.
@@
Financial Language Models: FinPLMs and FinLLMs build upon backbone models like BERT (Devlin et al., 2018) and LLaMa (Touvron et al., 2023) with large-scale financial texts. FinBERT (Araci, 2019) was the first FinPLM with open released corpus, later improved FinBERT (Yang et al., 2020) incorporating larger communication datasets. Models like FLANG Shah et al. (2022) combine BERT and ELECTRA (Clark et al., 2020). Chinese-specific PLMs like Mengzi (Zhang et al., 2021) and FinT5 (Lu et al., 2023) address unique financial tasks. BloombergGPT introduce the outstanding FinLLM, along with successors like FinGPT (Yang et al., 2023b), DISC-FinLLM (Chen et al., 2023), CFGPT (Li et al., 2023a), and FinMA (Xie et al., 2023b) excelling in specific data collection and supervised fine-tuning using backbone models Baichuan (Yang et al., 2023a) and InternLM (Team, 2023). However, monolingual FinLLMs struggle to provide unified cross-lingual finance services.
@@
Financial Evaluation Benchmarks: The first benchmark FLUE (Shah et al., 2022) and its successor FLARE (Xie et al., 2023b) evaluate classification, boundary detection, and question answering abilities in English domain. FLARE also evaluates time-series reasoning like forecasting stock price movements. BBT-CFLEB (Lu et al., 2023) is the first Chinese benchmark with 5 financial NLP tasks, while FinCUGE (Chen et al., 2023) focuses on 9 financial NLP tasks on social media and short news. FinEval (Zhang et al., 2023b) assesses LLMs through multiple-choice questions on finance, economy, and accounting. CFBenchmark (Lei et al., 2023) evaluates text processing abilities in recognition, classification, and generation, encompassing totaling 8 tasks, and CGCE (Zhang et al., 2023c) includes 200 general and 150 specific professional diverse QA tasks. However, most benchmarks lack comprehensive real-world application support, posing challenges for assessing bilingual FinLLMs.
3 Method
The overall framework of ICE-PIXIU, displayed in Figure 1, offers tailored services to various financial groups like banks, funds, and insurance firms. It offers professional solutions like event analysis, Q&A, stock forecasting, exams, credit ratings. ICE-PIXIU’s robust analytical prowess stems from meticulously curated 5 data types—Data Label from Classification (DLC), Data Label from Extraction (DLE), Data Text by Translation (DTT), Data Text with English (DTE), and Data Out of Field (DOF). These categories support various NLP tasks, encompassing classification (CLS), extraction (EXT), reasoning (REA), prediction (PRE), generation (GEN) and Translation (TRA), across 15 bilingual specific financial tasks in Chinese and English. Moreover, ICE-PIXIU not only boasts bilingual proficiency but also its broad support for multi-modal features, such as tables in financial reports and time series data in stock predictions.
3.1 Raw Data Construction
In this section, we introduce our bilingual Chinese-English raw data, designed for instruction tuning and evaluation benchmark. Compared with similar works, (Chen et al., 2023) focuses on CLS and REA financial NLP task, (Lei et al., 2023; Li et al., 2023a) emphasizes CLS and EXT financial NLP task. Our raw data comprises 40 different datasets covering almost all financial NLP tasks, including CLS, EXT, REA, PRE, GEN, and TRA. Unlike commonly self-instruct generation data methods (Wang et al., 2022), our data is sourced from open datasets in real-world financial scenarios. The sunburst chart illustrating the distribution of ICE-PIXIU across different levels is shown in Figure 2. As depicted, FinED and FinSM tasks boast datasets on a larger scale. Furthermore, Chinese data surpasses English data in terms of proportions in data types, NLP tasks, and specific tasks, particularly showcasing a more diverse distribution in specific financial tasks, covering a wider range of financial application scenarios. This indicates that compared to the English-focused PIXIU (Xie et al., 2023b), ICE-PIXIU demonstrates superior understanding of Chinese financial tasks. More details information of our raw data are in Appendix Table 7. According to statistics, the distribution percentages for all data types are: DLE (59.38%), DLC (20.08%), DTT (14.24%), DTE (5.96%), and DOF (0.34%), while the distribution percentages for Chinese and English NLP tasks are: ZH-CLS (18.83%), ZH-PRE (1.25%), ZH-EXT (56.64%), ZH-GEN (2.73%), ZH-TRA (14.24%), EN-CLS (1.61%), EN-PRE (3.31%), EN-EXT (0.05%), EN-REA (0.99%), and EN-DOF (0.34%).
3.2 Bilingual Financial Datasets
The ICE framework’s proficiency in bilingualism is benefited from the following Chinese financial data, English financial data, and Translation data.
3.2.1 Chinese Financial Datasets

### v3

@@
Abstract.
The advancement of Large Language Models (LLMs) has significantly enhanced financial analysis, but their use has been mostly limited to single languages, leaving the potential for bilingual Chinese-English capabilities unexplored. To bridge this gap, we introduce ICE-PIXIU111https://github.com/YY0649/ICE-PIXIU, the first open-source framework that establishes instruction datasets ICE-FIND, fine-tuned LLMs ICE-INTENT, and an evaluation benchmark ICE-FLARE for financial LLMs in both Chinese and English. ICE-FIND includes 18 Chinese-English bilingual specific tasks with 36 datasets and 604k data samples, expanding the breadth and depth of bilingual financial modeling. Harnessing this, we then develop the pioneering Chinese-English bilingual financial LLM ICE-INTENT by fine-tuning the InternLM backbone model. We evaluate our models and existing LLMs using ICE-FLARE, the first comprehensive bilingual evaluation benchmark with 40 datasets covering 10 NLP tasks and 20 bilingual specific tasks. Experimental results demonstrate significant multilingual performance disparities and biases in existing LLMs, with fine-tuning using translation data can enhance LLM’s cross-lingual generalization capabilities. ICE-INTENT models outperform SOTA LLMs like GPT-4 in Chinese financial tasks, attributed to diverse linguistic resources, high-quality human-annotated prompts, and strategic instruction fine-tuning.
@@ Abstract.
1. Introduction
As the influence of Chinese investments and consumption grows and English remains the dominant language in global business, effective cross-lingual communication between these languages is increasingly vital (Mohamed and Lashine, 2003). Given the close interaction between the financial markets of China and the United States, bridging linguistic and cultural gaps in the complex financial sector requires specialized tools and models tailored for cross-language business scenarios (Zhang et al., 2024; Nie et al., 2024a).
@@ 1. Introduction
Inspired by the success of general domain large language models (LLMs), the exploration of financial LLMs as financial cross-language solutions is gradually gaining popularity (Lu et al., 2023; Xie et al., 2023). Starting with the broad capabilities of models like OpenAI’s ChatGPT (Brown et al., 2020), the focus has shifted towards more specialized models such as BloombergGPT (Wu et al., 2023), which tackles the complex nuances of financial language and concepts. Subsequently, models like FinGPT (Yang et al., 2023a), InvestLM (Yang et al., 2023b) and FinTral (Bhatia et al., 2024) emerged, concentrating on English financial tasks. Notably, FinMA (Xie et al., 2023) addresses the challenges of instruction data diversity, marking a significant advancement.
@@
Beyond the efforts in English-centric models, Chinese financial LLMs (FinLLMs) have also made notable advances. Baidu’s XuanYuan (轩辕) (Zhang et al., 2023d) exemplifies this trend, focusing on the intricacies of Chinese financial language. Subsequent models such as DISC-FinLLM (Chen et al., 2023) and CFGPT (Li et al., 2023a) further enhance capabilities in financial analysis and reasoning. Recently, PanGu-π (Wang et al., 2023a) enhances a pre-trained model for better performance in financial examinations.
Although these advancements, there are significant challenges in their application and evaluation in bilingual contexts, especially regarding model development, diversity of instruction datasets, and evaluation methods in English and Chinese. As depicted in Table 1, there remain several issues: (1) All models are tailored exclusively for either English or Chinese, highlighting a significant lack of models that are proficient in both languages; (2) Moreover, most models are not open source, and particularly, nearly all crucial fine-tuning instruction datasets are closed, which will severely impede research in this field. (3) Current evaluation methods often focus on monolingual assessments and fail to comprehensively evaluate bilingual proficiency. Even more concerning is their lack of performance validation with data-out-of-training tasks, leading to potentially inflated evaluation metrics. (4) While existing instruction datasets including financial NLP tasks like classification, extraction, prediction and reasoning, these datasets, particularly for Chinese, show a notable lack of diversity in each task category. (5) Original and corresponding translated datasets are often overlooked in the development of comprehensive bilingual instruction datasets, making it challenging to assess the models’ cross-lingual generalisation.
@@
| FinLLMs | Backbone | Parameter Size | Language | Instruction | Evaluation | Open Source | Release Date | ||||
| zh | en | zh | en | dot | Model | Data | |||||
| BloombergGPT (Wu et al., 2023) | BLOOM (Le Scao et al., 2023) | 50B | en | 0 | 0 | 0 | 0 | 5 | 03/30/23 | ||
| XuanYuan2.0 (Zhang et al., 2023d) | BLOOM (Le Scao et al., 2023) | 176B | zh | - | 0 | 6 | 0 | 0 | 05/19/23 | ||
| FinMA (Xie et al., 2023) | LLaMA | 7/13B | en | 0 | 5 | 0 | 5 | 0 | 06/08/23 | ||
| FinGPT (Yang et al., 2023a) | ChatGLM (Du et al., 2021) | 7/13B | en | 0 | 6 | 0 | 6 | 0 | 06/09/23 | ||
| InvestLM (Yang et al., 2023b) | LLaMA (Touvron et al., 2023) | 65B | en | 0 | 4 | 0 | 4 | 0 | 09/15/23 | ||
| CFGPT (Li et al., 2023a) | InternLM (Team, 2023) | 7B | zh | 6 | 0 | 6 | 0 | 0 | 09/19/23 | ||
| DISC-FinLLM (Chen et al., 2023) | Baichuan (Yang et al., 2023c) | 13B | zh | 8 | 0 | 8 | 0 | 0 | 10/23/23 | ||
| PanGu-π (Wang et al., 2023a) | - | 1/7B | zh | - | 0 | 4 | 0 | 0 | 12/27/23 | ||
| FinTral (Bhatia et al., 2024) | Mistral (Jiang et al., 2023) | 7B | en | 0 | 9 | 0 | 9 | 0 | 02/16/24 | ||
| ICE-INTENT | InternLM (Team, 2023) | 7B | zh, en | 12 | 6 | 12 | 6 | 3 | 03/10/24 | ||
@@
To tackle this issue, we introduce ICE-PIXIU, the first open-source Chinese-English bilingual framework meticulously designed to bridge the research gap. As detailed in Table 2, ICE-PIXIU has painstakingly curated 40 datasets, incorporating 1,185,076 raw data points, 603,940 fine-tuning instruction data, and 95,091 evaluation data, strategically covering a diverse range of Chinese-English bilingual financial tasks. This diverse array not only underscores our model’s commitment to linguistic breadth but also to task-specific depth, ensuring ICE-INTENT’s proficiency across various financial scenarios 222Examples of each task are depicted in Appendix Table 6 to Table 8. ICE-PIXIU includes the bilingual open-source multi-task and Chinese-English instructional dataset (ICE-FIND) with 36 datasets spanning 18 specific tasks, the pioneering open-source Chinese-English bilingual evaluation benchmark (ICE-FLARE) with 40 datasets covering 20 specific tasks, and the open-source Chinese-English bilingual financial LLM, ICE-INTENT.
To establish a multi-task and bilingual instruction data and evaluation resources, we integrate Chinese and English financial specific tasks from 4 data categories, including (A) Chinese financial datasets, (B) English financial datasets, (C) Chinese translation datasets, and (D) out-of-training financial datasets (as outlined in Section 3.1.1), merging expert-annotated prompts with task-specific data samples. This led to the creation of Chinese and English bilingual financial instruction data called ICE-FIND. To enhance bilingual financial analysis capabilities, we introduce the bilingual financial LLM model, ICE-INTENT, achieved by fine-tuning the InternLM-7B backbone model with ICE-FIND. For evaluation, we further construct the bilingual ICE-FLARE benchmark by including test set from ICE-FIND, and introducing the data-out-of-training (DOT) task along with 3 unseen specific tasks and 4 unseen datasets. The inclusion of DOT task in the benchmark aims to assess the model’s generalization capabilities across various financial contexts and tasks.
We assess ICE-INTENT and baseline LLMs with the ICE-FLARE bilingual benchmark. Our analysis revealed three key findings: 1) Current LLMs, such as GPT-4, show a notable performance gap in Chinese financial tasks, highlighting disparities across languages. 2) The ICE-INTENT models excel in bilingual financial analysis, outperforming established models like GPT-4 in Chinese financial tasks. This success emphasizes the pivotal role of fine-tuning instructions and utilizing target language and high-resource datasets to enhance model capabilities. 3) Of particular note, fine-tuning LLMs with translated data not only bridges cross-lingual financial tasks but also significantly boosts the model’s performance in English tasks, demonstrating the beneficial effects of cross-lingual data transfer.
Our contributions can be summarized as follows: 1) Bilingual Capability Proficiency: ICE-INTENT, a bilingual financial LLM, showcases exceptional bilingual proficiency in English and Chinese, essential for navigating cross-lingual financial Q&A scenarios. 2) Diverse Chinese Financial Datasets: ICE-PIXIU enhances training and performance by integrating the Chinese classification, extraction, reasoning, and prediction NLP tasks, surpassing existing FinLLMs in task diversity and data scale. 3) Professionally Annotated Prompts: ICE-PIXIU provides a diverse, high-quality, expert-annotated set of prompts applied with similar fine-tuning instructions to enhance understanding of financial tasks. 4) Integration of English-to-Chinese Translation Data: The framework extends its capabilities by incorporating the English-to-Chinese translation datasets, bolstering bilingual training and cross-lingual generalization. 5) Cross-Lingual Evaluation with ICE-FLARE: ICE-PIXIU introduces ICE-FLARE, a rigorous cross-lingual evaluation benchmark that ensures consistent model performance in diverse linguistic contexts. 6) Open-Source Contribution: With its open-access approach, ICE-PIXIU offers its models, data, and evaluation resources to the research community, fostering collaborative advancement in financial NLP.
2. Related Work
@@
Financial Dataset Resource Financial pre-trained language models (FinPLMs) previously focus on single-task datasets but now most FinLLMs prioritize financial specific-task data collection. For instance, BloombergGPT (Wu et al., 2023) integrates general and financial text to train large 50B LLMs. FinLLMs such as FinGPT (Yang et al., 2023a; Zhang et al., 2023c) curate English instruction datasets for classification, extraction, and generation NLP tasks. PIXIU (Xie et al., 2023) provide open-source, multi-task and multi-modal instruction data covering various financial scenarios. BBT-Fin (Lu et al., 2023) Corpus aggregates Chinese financial data including corporate announcements, research reports, social media, and more, while XuanYuan 2.0 (Zhang et al., 2023d) provides a hybrid tuning dataset. Cornucopia’s (Yu, 2023) 26M instruction dataset covers application areas such as insurance and stocks. Most fine-tuning datasets are tailored for monolingual tasks, which limits their coverage.
@@
Financial Language Models FinPLMs and FinLLMs build on backbone models like BERT (Devlin et al., 2018) and LLaMa (Touvron et al., 2023) with large-scale financial texts. FinBERT (Araci, 2019) was the first FinPLM with an openly released corpus, later improved FinBERT (Yang et al., 2020) incorporating larger communication datasets. Models like FLANG (Shah et al., 2022) combine BERT and ELECTRA (Clark et al., 2020). Chinese-specific PLMs like Mengzi (Zhang et al., 2021) and FinT5 (Lu et al., 2023) address unique financial tasks. BloombergGPT introduces the outstanding FinLLM, along with successors like FinGPT (Yang et al., 2023a) and FinMA (Xie et al., 2023) excelling in specific data collection and supervised fine-tuning using backbone models Baichuan (Yang et al., 2023c) and InternLM (Team, 2023). However, monolingual FinLLMs struggle to offer consistent cross-lingual finance services.
@@
Financial Evaluation Benchmarks The first benchmark assesses classification, boundary detection, and question-answering in English, while FLARE (Xie et al., 2023) extends this to include time-series reasoning. BBT-CFLEB (Lu et al., 2023) is the first Chinese benchmark with 5 financial NLP tasks, while FinCUGE (Chen et al., 2023) focuses on 9 financial NLP tasks on social media and short news. FinEval (Zhang et al., 2023a) assesses LLMs through multiple-choice questions on finance, economics, and accounting. CFBenchmark (Lei et al., 2023) evaluates text processing abilities in recognition, classification, and generation, encompassing totaling 8 tasks, and CGCE (Zhang et al., 2023b) includes 200 general and 150 specific professionally diverse QA tasks. FinBen (Xie et al., 2024) serves as an extensive benchmark with 36 datasets over 24 financial tasks. However, these benchmarks lack support for cross-lingual and out-of-domain tasks, posing a challenge for evaluating the generalization ability of FinLLMs.
3. Method
3.1. ICE-FIND: Bilingual Financial Instruction Dataset
We present ICE-FIND, a Chinese-English bilingual financial instruction dataset, detailing the raw data sources and the instruction construction process using high-quality prompts by human evaluation. Unlike current resources, ICE-FIND emerges as the first instruction data tailored for bilingual FinLLMs, showcasing originality.
3.1.1. Raw Data Integration
ICE-FIND is built from open-source data, covering a range of specific tasks in real-world financial scenarios. It stands out by utilizing expert-curated, high-quality sources, offering cost-efficiency over self-instruct data generation methods (Wang et al., 2022) without usage restrictions, and providing varied data formats and multilingual support for flexible task adaptation. As shown in Table 2, it curated 5 data types-Data-Label-from-Classification (DLC), Data Labels from-Extraction (DLE), Data-Text-by-Translation (DTT), Data-Text-with-English (DTE), and Data-Out-of-Training (DOT). These categories support 10 financial NLP tasks such as Classification (CLS), Extraction (EXT), Reasoning (REA), Prediction (PRE), Generation (GEN), and Translation (TRA) across 12 Chinese and 6 English specific tasks. Compared with similar studies (Wang et al., 2023b; Chen et al., 2023) respectively focus on PRE, as well as CLS and REA NLP task, (Lei et al., 2023; Li et al., 2023a) emphasizes CLS and EXT NLP task. Our raw data comprises 40 different datasets covering almost all financial NLP tasks, including CLS, EXT, REA, PRE, GEN, and TRA.
Our proposed ICE-FIND consists of four datasets as follows: (A) Chinese financial datasets, (B) English financial datasets, (C) Chinese translation datasets, and (D) out-of-training financial datasets.
| Language | Data Type | NLP Task | Specific Task | Dataset | Raw | Instruction | Evaluation | Data Source | License |
| StockB | 9,812 | 9,812 | 1,962 | social texts | Apache-2.0 | ||||
| FinSM | BQC | 120,000 | 110,000 | 10,000 | bank service logs | Public | |||
| AFQMC | 38,650 | 38,650 | 4,316 | online chat service | Apache-2.0 | ||||
| FinNC | NL | 7,955 | 7,955 | 884 | news articles | Public | |||
| NL2 | 7,955 | 7,955 | 884 | news articles | Public | ||||
| FinNJ | NSP | 4,499 | 4,499 | 500 | social texts | Public | |||
| FinAS | FinevalF | 1,115 | 1,115 | 222 | financial exam | Apache-2.0 | |||
| FinRE | RE | 14,973 | 14,973 | 1,489 | news, entity pairs | Public | |||
| FinER | CNER | 1,685 | 1,685 | 337 | financial reports | Public | |||
| 20CCKS | 372,810 | 45,796 | 9,159 | news, reports | CC BY-SA 4.0 | ||||
| 21CCKS | 8,000 | 7,000 | 1,400 | news, reports | CC BY-SA 4.0 | ||||
| 22CCKS | 109,555 | 59,143 | 11,829 | news, reports | CC BY-SA 4.0 | ||||
| FinQA | CEnQA | 8,281 | 668 | 133 | earnings reports | MIT license | |||
| CConFinQA | 12,594 | 1,189 | 237 | earnings reports | MIT license | ||||
| FiQA-SA | 1,173 | 1,173 | 235 | news headlines,tweets | Public | ||||
| FinCC | German | 1,000 | 1,000 | 200 | credit records | CC BY-SA 4.0 | |||
| Australian | 690 | 690 | 139 | credit records | CC BY-SA 4.0 | ||||
| ConFinQA | 3,458 | 12,594 | 1,490 | earnings reports | MIT License | ||||
| DOT | EN-DOT | FinER | FINER-ORD | 1,075 | - | 1,075 | news articles | CC BY-SA 4.0 | |
| FinTS | ECTSUM | 495 | - | 495 | earning call transcipts | Public | |||
| EDTSUM | 2,000 | - | 2,000 | news articles | Public | ||||
| FinDC | FOMC | 496 | - | 496 | FOMC transcripts | CC BY-SA 4.0 |